OpenStudy (anonymous):

Evaluate each of the following (possibly infinite) limits or explain why it does not exist lim x→5− lnx/(x−5)^2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5-} \ln x/(x-5)^2\]

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 5-} {\ln x \over (x-5)^2} = {\ln 5 \over 0^2} = \infty\]

OpenStudy (anonymous):

i think the denominator cant be 0 right ?

OpenStudy (anonymous):

Yes ... you are right!! Also try to graph the function \[ {\ln x \over (x-5)^2}\] at x -> 5, the graph goes up to infinity ( where it diverges)

OpenStudy (anonymous):

can u show me some working how to do it ?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you are welcome :D